OpenStudy (vortish):

find the slope of the line that passes thru points (1980,16000) and (2000,43000)

OpenStudy (anonymous):

m(slope)= change in y/change in x m=(43000-16000)/(2000-1980), dont have a calculator so slove it your self. Note i have assumed that (2000,43000) is the second point.

OpenStudy (diyadiya):

Slope of the line passing through the points (x1,y1) & (x2,y2) is given by the formula \[\Large slope = \frac{y2-y1}{x2-x1}\]

OpenStudy (diyadiya):

Here (x1,y1) = (1980,16000) (x2,y2) = (2000,43000) that means x1 = 1980 y1 = 16000 x2 = 2000 y2 = 43000 Now substitute it to the formula \[\large Slope = \frac{y2-y2}{x2-x1}\] posideon is completely right!